# Computing finite–temperature elastic constants with noise cancellation

### v1

@@ V Conclusions
We validated the method across a diverse set of material systems, beginning with crystalline argon as a simple benchmark and progressing through ordered and amorphous silicon to structurally complex amorphous polymers such as poly(methyl methacrylate) and cellulose derivatives. In each case, our results show rapid and stable convergence of the elastic constants, with reasonable agreement with available reference data. For the polymeric systems, especially cellulose and cellulose acetate, our simulations provide the first calculation of the full elastic tensor at finite temperature, offering valuable new data for the materials modeling community. We have further applied this method to determine the temperature–dependent elastic constants of a cellulose derivative and their direct implications for thermal transport [29], as well as the elastic constants of experimentally relevant commodity polymers under high pressures [39].
@@
Acknowledgment: D.M. thanks Tiago Espinosa de Oliveira for a collaboration [29], which provided key systems investigated in this work. Financial support for D.M. and M.M. was provided by the Bundesministerium für Technologie, Forschung und Raumfahrt (BMTRF) within the project 16ME0658K MExMeMo and European Union– NextGenerationEU. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss–centre.eu) for providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC) where the polymer simulations were performed.

### v2

@@ V Conclusions
We validated the method across a diverse set of material systems, beginning with crystalline argon as a simple benchmark and progressing through ordered and amorphous silicon to structurally complex amorphous polymers such as poly(methyl methacrylate) and cellulose derivatives. In each case, our results show rapid and stable convergence of the elastic constants, with reasonable agreement with available reference data. For the polymeric systems, especially cellulose and cellulose acetate, our simulations provide the first calculation of the full elastic tensor at finite temperature, offering valuable new data for the materials modeling community. We have further applied this method to determine the temperature–dependent elastic constants of a cellulose derivative and their direct implications for thermal transport [36], as well as the elastic constants of experimentally relevant commodity polymers under high pressures [49] and polyelectrolyte melts [50]. While the current work utilizes a simple forward-difference scheme to demonstrate the robust performance of the noise cancellation technique even under non-ideal conditions, moving to a central-difference scheme would offer an immediate path to further increasing numerical stability. Such an approach would likely simplify the optimization of strain magnitudes by extending the range over which anharmonic effects remain negligible.
The method proposed in this work targets exclusively the (quasi–)static elastic constants. The noise–cancellation technique still bears the possibility to deduce viscoelastic response functions. To this end, a pair of constant–stress simulations would have to be performed, one under zero stress and another one under a constant perturbative stress. The (expectation value of the) difference in the strain between these two runs, divided by the stress increment, allows the time–dependent and thus frequency–dependent elastic modulus to be deduced. The signal–to–noise ratio in this approach will certainly be much improved compared to a direct measurement without noise cancellation. Nonetheless, decorrelation of the signal, which is expected to set in once the long–wavelength modes start relaxing, may require special measures to ensure meaningful results at medium and large time scales.
@@
Acknowledgment: D.M. thanks Tiago Espinosa de Oliveira for a collaboration [36], which provided key systems investigated in this work. Financial support for D.M. and M.M. was provided by the Bundesministerium für Technologie, Forschung und Raumfahrt (BMTRF) within the project 16ME0658K MExMeMo and European Union– NextGenerationEU. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss–centre.eu) for providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC) where the polymer simulations were performed. M.H.M. acknowledges helpful discussions with Sergey Suhomlinov and Lars Pastewka.